# JARVIS-VLA: Post-Training Large-Scale Vision Language Models to Play Visual Games with Keyboards and Mouse

### v1

@@ 3.2 VLA Performance Evaluation
We present the performance results of our proposed model across four categories from the MCU benchmark [31], as shown in Table 1. For each MCU task, we collect over 100 random trajectories, which are used to fine-tune base VLMs to create our final VLA models.
@@
We evaluate three variants of the VLMs as base models: 1) Qwen2-VL (raw): the original VLM checkpoint fine-tuned for downstream tasks. 2) Qwen2-VL (IL): post-trained on large-scale cross-task trajectories (distinct from downstream tasks) using imitation learning. 3) JARVIS-VLA-Qwen2: post-trained on our proposed off-trajectory vision-language tasks. Performance is measured by the average success rate across tasks within each category. Our results show that JARVIS-VLA-Qwen2-VL, post-trained using our approach, consistently outperforms prior methods across almost all tasks.
@@
Remarkably, even without task-specific post-training, raw Qwen2-VL model, fine-tuned on downstream tasks, outperforms several previous baselines, including STEVE-1 [30] and GROOT [12], which were trained using large-scale imitation learning. This highlights the effectiveness of using a robust pre-trained VLM as the base model for the policy, leading to strong performance even without additional fine-tuning.
@@
Notably, we observe a significant performance boost with ActVLP post-training. For tasks such as Craft Items and Smelt Items, where previous methods struggled, JARVIS-VLA-Qwen2-VL achieves success rates more than double those of the baseline models. This underscores the effectiveness of our off-trajectory vision-language task strategy. Furthermore, JARVIS-VLA-Qwen2-VL outperforms Qwen2-VL (IL) by over 15%, despite using only 21% of the training trajectory data. In crafting category tasks, the JARVIS-VLA model surpasses traditional baselines by more than double, outperforming models like VPT-BC [4] and STEVE-1 [30] on tasks such as "Craft crafting table" ([Uncaptioned image]). This significant improvement is primarily due to the use of ViT in VLM and high-resolution processing, which are crucial for tasks like crafting and smelting that demand precise control in the GUI interface. This suggests that integrating off-trajectory vision-language tasks into the training pipeline enhances decision-making capabilities, enabling more accurate action predictions in VLA models. Further analysis and additional experiments will be presented in the next section.
@@ 3.3 Ablation on Non-Trajectory Datasets
To understand the impact of different task enhancements, we conduct an ablation study by dividing the non-trajectory datasets and training Qwen2-VL separately on three types of tasks: spatial grounding, visual language alignment, and knowledge-based question-answering, which are all related to Minecraft games. This results in three variants of the VLM, each augmented with one of these capabilities—spatial grounding, visual recognition, and world knowledge. All models are finetuned using the same gameplay dataset and imitation learning techniques. We also develop a benchmark, detailed in Appendix E, to evaluate these capabilities. For this evaluation, we select three long-sequence atomic tasks: "Craft the diamond sword" ([Uncaptioned image]), "Mine the obsidian" ([Uncaptioned image]), and "Cook the beef" ([Uncaptioned image]), as downstream instruction-following tasks.
The results of our ablation studies, presented in Figure 4, demonstrate that post-training with non-trajectory vision-language tasks significantly enhances the core capabilities of the VLM across the respective benchmarks. Notably, after fine-tuning, models enhanced with spatial grounding exhibit the most substantial improvement in downstream decision-making tasks. These findings underscore the effectiveness of non-trajectory post-training in boosting the performance of Vision-Language-Action models in decision-making tasks, even when the focus is on a single task. We find that non-trajectory vision-language tasks, which are essential for agent pipelines [44, 45], are more effective for fine-tuning end-to-end VLA models. This demonstrates the connection between developing LLM-based agent pipelines with separate modules and fine-tuning end-to-end VLA models.
@@ 3.4 Scaling Experiments
Recent work on large language models (LLMs) trained on vast amounts of text via next-token prediction has shown strong scaling laws [48, 19, 32, 47]. We investigate whether VLAs, obtained through post-training on VLMs, exhibit similar scaling behavior. Specifically, we explore two questions: Q1) Can scaling up downstream imitation learning trajectories further improve the VLA’s task success rate? Q2) Does increasing the amount of non-trajectory vision-language tasks used during post-training enhance task completion success?
@@
The results for Q2 are illuminated in Figure 6. We also explore the relationship between the evaluation loss during post-training on non-trajectory vision-language tasks and task success rate in downstream tasks. We use base models from different stages of post-training (with different eval loss on post-training datasets), fine-tuning them with the same downstream trajectory dataset. The baseline represents post-training using imitation learning on cross-task trajectories. We find that, for nearly all tasks, the success rate in downstream tasks correlates linearly with evaluation loss in post-training, with the lowest loss yielding the best results. Notably, models post-trained with knowledge-based tasks exhibit the best downstream performance for a given evaluation loss. Models enhanced with spatial grounding show the lowest evaluation loss and the highest task success rates. These findings demonstrate that scaling up off-trajectory vision-language datasets directly enhances downstream task performance, which has been overlooked in previous VLA works [4, 25, 5].
@@ 4.1 Visual-Language-Action Models
Imitation learning (IL) involves learning by mimicking expert interactions with the environment, with the primary challenge being the collection of high-quality expert demonstration datasets. Numerous studies have sought to enhance traditional IL approaches [15, 39, 6]. A promising direction is the use of Visual-Language-Action (VLA) models [7, 25, 55, 51, 57], which adopt end-to-end imitation learning by fine-tuning VLMs. OpenVLA [25] has demonstrated the importance of selecting a capable VLM backbone, a conclusion further reinforced by RoboVLM [28]. Similarly, Brohan et al. [7] highlighted that co-training with web-scale vision-language data significantly improves the generalization of VLA models. While previous works primarily focused on optimizing the selection of VLMs, several recent studies have begun to pay attention to the comprehension capabilities of VLA models [52, 60, 59, 13]. However, few have explicitly focused on enhancing the VLM backbone itself through visual-language post-training. Our work addresses this gap by proposing targeted visual-language post-training methods to enrich the capabilities of VLMs, thereby improving their performance on downstream VLA tasks.
@@ 4.2 VLM-based Agents in Minecraft
Existing Minecraft agents based on VLMs typically adopt hierarchical architectures [44, 58, 21, 17, 53, 9]. These methods leverage a VLM’s world knowledge for planning via zero-shot or few-shot in-context learning, without modifying the VLM parameters during agent optimization [44, 45, 42, 29]. STEVE-EYE [56] fine-tuned Llama language models [40] using internet text data, achieving improved planning over zero-shot prompting. MineDreamer [58] employs the instruction-following capability of VLMs to predict future visual observations and generate actions based on STEVE-1 [30]. OmniJARVIS [46] uses a behavior tokenizer [12, 11] to model human trajectories in Minecraft with pretrained VLMs. While these approaches optimize VLMs, they still rely on additional policies for action grounding. In contrast, we propose a VLA-based agent model that generates actions directly from textual instructions and visual inputs, eliminating the need for extra grounding policies.
@@ 5 Conclusions
Looking ahead, there are several avenues for improvement in future work. First, it is crucial to enhance the inference throughput of JARVIS-VLA, which is currently constrained by the large parameter size of the VLA based on VLM [8]. We believe that future integration with MoE [24, 22] could further improve the model’s inference efficiency, with the goal of achieving gameplay performance levels exceeding 40Hz. Additionally, there remains potential for further performance gains. While JARVIS-VLA outperforms previous Minecraft policies, it still falls short of the performance demonstrated by top human players, who achieve success rates above 90%.
@@ Appendix A Observation and Action Space
We rely solely on visual images for observation, without any symbolic information, similar to VPT [4]. To closely resemble the actions of real human players, our action space covers all possible player actions excluding the typing of arbitrary letters. For the keypress and click actions, we allocate corresponding reserved tokens from outside the original VLM tokenizer vocabulary. Regarding mouse movements, we largely follow the approach adopted by VPT [4], applying the mu-law encoding method to discretize mouse X and Y actions separately into 21 bins, totaling 42 discrete bins. These bins are similarly mapped to reserved tokens. It is worth mentioning that although Qwen2-VL [43] does not explicitly provide reserved tokens like Llama3 [33], it is still feasible to achieve equivalent results by expanding the number of special tokens, due to the fact that the size of the vocabulary is smaller than the dimension of the word embeddings.
During inference, models generate actions in a token-by-token manner—first predicting the keybuttons to be pressed, followed by the camera Y and camera X values.
@@ Appendix B Training Configurations
The training configurations for both Visual-Language Post-Training and Action Post-Training are largely consistent. All experiment were conducted on NVIDIA A800-SXM4-80GB GPUs, utilizing CUDA version 12.1 and Hugging Face Transformers version 4.47.0. Both training stages utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay were set to 0, and $\epsilon=1\times 10^{-8}$. A cosine learning rate schedule was adopted with the learning rate of $5\times 10^{-6}$ and a warmup of 200 steps. The training used bfloat16 precision, a maximum gradient norm of 1.0, and a fixed random seed of 42. To accelerate training, DeepSpeed with ZeRO-1 [36]optimization was employed. For Visual-Language Post-Training, the maximum token length was set to 3584, and we set a batch size per device of 2 and a gradient accumulation of 4 . For Action Post-Training, the maximum token length was set to 512, which allowed a batch size per device of 8 and a gradient accumulation of 1 step per update. Ensuring that the total batch size remained 256. Both stages were trained using 32 A800 GPUs, with the Visual-Language Post-Training phase running for 128 GPU hours and the Action Post-Training phase running for 512 GPU hours.
@@ Appendix C Details of Inference
During inference, we require JARVIS-VLA to output in the following format which is as same as the gameplay dataset:
Example of JARVIS-VLA Interaction for One Turn of Iteration Instruction: Craft a bread so I can use it. Arrange the materials in the crafting grid according to the following pattern: wheat | wheat | wheat wheat | wheat | wheat and get 1 bread. Observation: [Uncaptioned image] Action: <|action_begin|><|cam_w_13|><|cam_h_5|><|action_end|>
To accelerate inference, we utilized vllm [26] to load JARVIS-VLA in bfloat16 precision on four A800 GPUs. The frames-per-second (FPS) reached 55.
Example of JARVIS-VLA Interaction for Multi Turn of Iteration Instruction: Kill a zombie. [Uncaptioned image] Instruction: craft a crafting table. [Uncaptioned image]

### v2

@@ 3.2 VLA Performance Evaluation
We present the performance results of our proposed model across four categories from the MCU benchmark (Lin et al., 2023), as shown in Table 1. For each MCU task, we collect over 100 random trajectories, which are used to fine-tune base VLMs to create our final VLA models.
@@
We evaluate three variants of the VLMs as base models: 1) Qwen2-VL (raw): the original VLM checkpoint, fine-tuned directly on the task-specific dataset. 2) Qwen2-VL (IL): first undergoes action post-training, then trained on the same task-specific fine-tuning dataset. 3) JARVIS-VLA-Qwen2-VL: first post-trained using our proposed ActVLP paradigm, then fine-tuned on the same dataset. Performance is measured by the average success rate across tasks within each category. Our results show that JARVIS-VLA-Qwen2-VL, post-trained using our approach, consistently outperforms prior methods across almost all tasks.
@@
Remarkably, even without task-specific post-training, raw Qwen2-VL model, fine-tuned on downstream tasks, outperforms several previous baselines, including STEVE-1 (Lifshitz et al., 2024) and GROOT (Cai et al., 2024c), which were trained using large-scale imitation learning. This highlights the effectiveness of using a robust pre-trained VLM as the base model for the policy, leading to strong performance even without additional fine-tuning.
@@
Notably, we observe a significant performance boost with ActVLP post-training. For tasks such as Craft Items and Smelt Items, where previous methods struggled, JARVIS-VLA-Qwen2-VL achieves success rates more than double those of the baseline models. This underscores the effectiveness of our off-trajectory vision-language task strategy. Furthermore, JARVIS-VLA-Qwen2-VL outperforms Qwen2-VL (IL) by over 15%, despite using only 21% of the training trajectory data. In crafting category tasks, the JARVIS-VLA model surpasses traditional baselines by more than double, outperforming models like VPT-BC (Baker et al., 2022) and STEVE-1 (Lifshitz et al., 2024) on tasks such as "Craft crafting table" ([Uncaptioned image]). This significant improvement is primarily due to the use of ViT in VLM and high-resolution processing, which are crucial for tasks like crafting and smelting that demand precise control in the GUI interface. This suggests that integrating off-trajectory vision-language tasks into the training pipeline enhances decision-making capabilities, enabling more accurate action predictions in VLA models. Further analysis and additional experiments will be presented in the next section.
3.3 Ablation on Training Paradigm
To assess the impact of visual-language post-training, we designed an additional baseline, Qwen2-VL-7B (one-stage). It is co-trained on the original checkpoint of Qwen2-VL, using the combination of non-trajectory datasets and trajectory datasets in a single unified training stage. After training for one epoch, Qwen2-VL-7B (one-stage) is fine-tuned on MCU tasks. We compare it with JARVIS-VLA-Qwen2-VL, which is post-trained in three separated steps.
Model [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Qwen2-VL(raw) 0.53 0.40 0.03 0.10
Qwen2-VL-7B (one-stage) 0.10 0.40 0.07 0.13
ActVLP-Qwen2-VL 0.83 0.63 0.77 0.70
Table 2: Ablation results for different training paradigms. We use Qwen2-VL (raw) as the baseline and compare it against Qwen2-VL-7B (one-stage) and JARVIS-VLA-Qwen2-VL on four Minecraft tasks: "craft diamond sword", "craft ladder", "cook beef", and "smelt iron ingot".
The results presented in Table 2 indicate that separating visual-language post-training from action learning, rather than merging all stages into a single training phase, yields a significant performance boost. For instance, when comparing our JARVIS-VLA-Qwen2-VL to the one-stage co-finetuning approach, we observed substantial improvements, including approximately 57% on the "Smelt iron ingot [Uncaptioned image]" task and 23% on the "Crafting the ladder [Uncaptioned image]" task. We hypothesize that dedicated post-training on visual-language datasets enhances the foundational capabilities of the VLM, providing a more advantageous starting point for subsequent imitation learning.
@@ 3.4 Ablation on Non-Trajectory Datasets
To understand the impact of different task enhancements, we conduct an ablation study by dividing the non-trajectory datasets and training Qwen2-VL separately on three types of tasks: spatial grounding, vision language alignment, and knowledge-based question-answering, which are all related to Minecraft games. This results in three variants of the VLM, each augmented with one of these capabilities—spatial grounding, visual recognition, and world knowledge. All models are finetuned using the same gameplay dataset and imitation learning techniques. We also develop a benchmark, detailed in Appendix E, to evaluate these capabilities. For this evaluation, we select three long-sequence atomic tasks: "Craft the diamond sword" ([Uncaptioned image]), "Mine the obsidian" ([Uncaptioned image]), and "Cook the beef" ([Uncaptioned image]), as downstream instruction-following tasks.
The results of our ablation studies, presented in Figure 4, demonstrate that post-training with non-trajectory vision-language tasks significantly enhances the core capabilities of the VLM across the respective benchmarks. Notably, after fine-tuning, models enhanced with spatial grounding exhibit the most substantial improvement in downstream decision-making tasks. These findings underscore the effectiveness of non-trajectory post-training in boosting the performance of Vision-Language-Action models in decision-making tasks, even when the focus is on a single task. We find that non-trajectory vision-language tasks, which are essential for agent pipelines (Wang et al., 2023, 2024c), are more effective for fine-tuning end-to-end VLA models. This demonstrates the connection between developing LLM-based agent pipelines with separate modules and fine-tuning end-to-end VLA models.
@@ 3.5 Scaling Experiments
Recent work on large language models (LLMs) trained on vast amounts of text via next-token prediction has shown strong scaling laws (Wei et al., 2022; Du et al., 2024; Lin et al., 2024; Wang et al., 2024e). We investigate whether VLAs, obtained through post-training on VLMs, exhibit similar scaling behavior. Specifically, we explore two questions: Q1) Can scaling up downstream imitation learning trajectories further improve the VLA’s task success rate? Q2) Does increasing the amount of non-trajectory vision-language tasks used during post-training enhance task completion success?
@@
The results for Q2 are illuminated in Figure 6. We also explore the relationship between the evaluation loss during post-training on non-trajectory vision-language tasks and task success rate in downstream tasks. We use base models from different stages of post-training (with different eval loss on post-training datasets), fine-tuning them with the same downstream trajectory dataset. The baseline represents post-training using imitation learning on cross-task trajectories. We find that, for nearly all tasks, the success rate in downstream tasks correlates linearly with evaluation loss in post-training, with the lowest loss yielding the best results. Notably, models post-trained with knowledge-based tasks exhibit the best downstream performance for a given evaluation loss. Models enhanced with spatial grounding show the lowest evaluation loss and the highest task success rates. These findings demonstrate scaling up off-trajectory vision language datasets directly enhances downstream task performance, which has been overlooked in previous VLA works (Kim et al., 2024; Brohan et al., 2023).
@@ 4.1 Visual-Language-Action Models
Imitation learning (IL) involves learning by mimicking expert interactions with the environment, with the primary challenge being the collection of high-quality expert demonstration datasets. Numerous studies have sought to enhance traditional IL approaches (Chi et al., 2023; Team et al., 2024; Brohan et al., 2022). A promising direction is the use of Visual-Language-Action (VLA) models (Brohan et al., 2023; Kim et al., 2024; Zheng et al., 2024; Zhang et al., 2024; Zhong et al., 2025), which adopt end-to-end imitation learning by fine-tuning VLMs. OpenVLA (Kim et al., 2024) has demonstrated the importance of selecting a capable VLM backbone, a conclusion further reinforced by RoboVLM (Li et al., 2024b). Similarly, Brohan et al. (2023) highlighted that co-training with web-scale vision-language data significantly improves the generalization of VLA models. While previous works primarily focused on optimizing the selection of VLMs, several recent studies have begun to pay attention to the comprehension capabilities of VLA models (Zhang et al., 2025; Zhu et al., 2025; Zhou et al., 2025; Chen et al., 2025). However, few have explicitly focused on enhancing the VLM backbone itself through visual-language post-training. Our work addresses this gap by proposing targeted visual-language post-training methods to enrich the capabilities of VLMs, thereby improving their performance on downstream VLA tasks.
@@ 4.2 VLM-based Agents in Minecraft
Existing Minecraft agents based on VLMs typically adopt hierarchical architectures (Wang et al., 2023; Zhou et al., 2024; Fan et al., 2022; Deng et al., 2025; Zhao et al., 2024; Cai et al., 2023). These methods leverage a VLM’s world knowledge for planning via zero-shot or few-shot in-context learning, without modifying the VLM parameters during agent optimization (Wang et al., 2023, 2024c, 2024a; Li et al., 2024c). STEVE-EYE (Zheng et al., 2023) fine-tuned Llama language models (Touvron et al., 2023) using internet text data, achieving improved planning over zero-shot prompting. MineDreamer (Zhou et al., 2024) employs the instruction-following capability of VLMs to predict future visual observations and generate actions based on STEVE-1 (Lifshitz et al., 2024). OmniJARVIS (Wang et al., 2024d) uses a behavior tokenizer (Cai et al., 2024c, b) to model human trajectories in Minecraft with pretrained VLMs. While these approaches optimize VLMs, they still rely on additional policies for action grounding. In contrast, we propose a VLA-based agent model that generates actions directly from textual instructions and visual inputs, eliminating the need for extra grounding policies.
@@ 5 Conclusions
Looking ahead, there are several avenues for improvement in future work. First, it is crucial to enhance the inference throughput of JARVIS-VLA, which is currently constrained by the large parameter size of the VLA based on VLM (Budzianowski et al., 2024). We believe that future integration with MoE (Jacobs et al., 1991; Fedus et al., 2022) could further improve the model’s inference efficiency, with the goal of achieving gameplay performance levels exceeding 40Hz. Additionally, there remains potential for further performance gains. While JARVIS-VLA outperforms previous Minecraft policies, it still falls short of the performance demonstrated by top human players, who achieve success rates above 90%.
@@ Appendix A Observation and Action Space
We rely solely on visual images for observation, without any symbolic information, similar to VPT (Baker et al., 2022). Notably, we increase the image resolution to 644×364, providing significantly more visual detail than VPT (128×128), which helps JARVIS-VLA better perceive the environment.
To closely mimic human behavior, our action space includes all possible player actions except for arbitrary text input. We allocate reserved tokens outside the original VLM vocabulary to represent keypress and click actions. For mouse movements, we follow VPT (Baker et al., 2022), using $\mu$-law encoding to discretize X and Y axes separately into 21 bins (42 tokens total), each mapped to a reserved token. Although Qwen2-VL (Wang et al., 2024b) does not explicitly support reserved tokens like Llama3 (Meta, 2024), we can add special tokens by expanding the vocabulary, since its embedding table is underutilized. Table 3 shows the action space JARVIS-VLA uses. During inference, the model predicts actions token by token: key/button actions first, followed by camera movements (Pitch and Yaw).
Table 3: The action space we use in Minecraft.
Index Action Human Action Description
1 Forward key W Move forward.
2 Back key S Move backward.
3 Left key A Strafe left.
4 Right key D Strafe right.
5 Jump key Space Jump. When swimming, keeps the player afloat.
6 Sneak key left Shift Slowly move in the current direction of movement.
7 Sprint key left Ctrl Move quickly in the direction of current motion.
8 Attack left Button Destroy blocks (hold down); Attack entity (click once).
9 Use right Button Interact with the block.
10 hotbar:[1-9] keys 1 - 9 Selects the appropriate hotbar item.
11 Inventory key E Open/Close the inventory.
12 Yaw move Mouse X Turning; aiming; camera movement.
13 Pitch move Mouse Y Turning; aiming; camera movement.
@@ Appendix B Training Configurations
The training configurations for both Visual-Language Post-Training and Action Post-Training are largely consistent. All experiment were conducted on NVIDIA A800-SXM4-80GB GPUs, utilizing CUDA version 12.1 and Hugging Face Transformers version 4.47.0. Both training stages utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay were set to 0, and $\epsilon=1\times 10^{-8}$. A cosine learning rate schedule was adopted with the learning rate of $5\times 10^{-6}$ and a warmup of 200 steps. The training used bfloat16 precision, a maximum gradient norm of 1.0, and a fixed random seed of 42. To accelerate training, DeepSpeed with ZeRO-1 (Rasley et al., 2020)optimization was employed. For Visual-Language Post-Training, the maximum token length was set to 3584, and we set a batch size per device of 2 and a gradient accumulation of 4 . For Action Post-Training, the maximum token length was set to 512, which allowed a batch size per device of 8 and a gradient accumulation of 1 step per update. Ensuring that the total batch size remained 256. Both stages were trained using 32 A800 GPUs, with the Visual-Language Post-Training phase running for 128 GPU hours and the Action Post-Training phase running for 512 GPU hours.
@@ Appendix C Details of Inference
During inference, JARVIS-VLA is expected to output actions in a format consistent with the gameplay dataset, as illustrated below:
Example of JARVIS-VLA Interaction for One Turn of Iteration Instruction: Craft a bread so I can use it. Arrange the materials in the crafting grid according to the following pattern: wheat | wheat | wheat wheat | wheat | wheat and get 1 bread. Observation: [Uncaptioned image] Action: <|action_begin|><|cam_w_13|><|cam_h_5|><|action_end|> <|action_begin|><|use|><|cam_w_3|><|cam_h_2|><|action_end|>
Action Chunking Strategy
To optimize both inference speed and predictive performance, we designed an adaptive action chunking strategy that varies the number of predicted actions per observation across different operational phases. The strategy begins with a chunk size of one action per observation during Action Post-Training to ensure precise policy alignment. In the subsequent fine-tuning stages, the chunk size is increased to three to accelerate learning. Finally, during inference, the model generates two actions per prediction step, balancing stability and efficiency.
To validate this multi-phase chunking approach, we conducted experiments on the MCU (Lin et al., 2023) tasks using Qwen2-VL-2B (Wang et al., 2024b) as the base model. The results are presented in  Table 4. This ablation study suggests that employing longer action chunks during advanced training stages and a moderate chunk size for inference can improve both success rates and FPS.
Table 4: Ablation study on action chunk sizes. "APT Chunks" denotes Action Post-Training chunks, and "FT Chunks" denotes Subsequent Fine-tuning chunks.
APT Chunks FT Chunks Inference Chunks FPS Success Rate
1 1 1 8 0.47
1 3 1 8 0.53
1 3 2 15 0.67
1 3 3 21 0.20
3 3 1 8 0.03
As shown in Table 4, the chunking strategy of using one action per observation during Action Post-Training, three during Further Fine-tuning, and two during Inference (1 $\rightarrow$ 3 $\rightarrow$ 2) achieved the highest success rate (0.67) and a solid frame rate (15 FPS) in this specific ablation. We did not explore chunk sizes larger than three, as the Minecraft environment changes rapidly with each discrete action. Predicting multiple future actions in such a dynamic setting poses a significant challenge.
It is also important to clarify the choice of inference chunk size in our main results. To ensure a rigorous and fair comparison with baseline models such as Qwen2-VL (raw) and Qwen2-VL (IL)—which both operate with a single predicted action—we configured JARVIS-VLA to use an inference chunk size of one in our primary experiments. While this setting facilitates direct comparison, our ablation results 4 suggest that a chunk size of two at inference may further improve both efficiency and task success under alternative evaluation criteria.
Infrastructure for Efficient Inference
To facilitate efficient inference, JARVIS-VLA was served using the vLLM (Kwon et al., 2023) in bfloat16 precision. This setup was typically deployed on four A800 GPUs. Compared to a standard Hugging Face Transformers implementation, vLLM increased the inference throughput of JARVIS-VLA by up to 1.7 times, without introducing any significant degradation in task performance.
Achieved Inference Performance
The final inference speed reflects both the action chunking strategy and infrastructure optimizations. In our main results, which use single-action inference for fair comparison, JARVIS-VLA-Qwen2-VL-7B runs at 5.5 FPS, while JARVIS-VLA-Qwen2-VL-2B achieves 8 FPS—consistent with the results in rows 1 and 2 of Table 4. Notably, JARVIS-VLA-Qwen2-VL-2B can reach 15 FPS with a chunk size of two, highlighting a trade-off we made in favor of direct benchmarking with baseline models.